OpenStudy (anonymous):

1. Find the area of a rhombus if one diagonal has length 30 and a side has length 17.

OpenStudy (anonymous):

need helpppppppppppppppppp!!!!!!!!!

OpenStudy (anonymous):

help me?

OpenStudy (azureilai):

ok. So first, I will explain how to do the problem. Let "d1" represent the first diagonal and "d2" represent the second diagonal. Also let "a" represent the side length. I'm going to solve this by using the diagonal method, which is (d1*d2)/2. You already have one diagonal, so we need to figure out the other diagonal. You can do that by using this formula. d1^2=d2^2=4a^2. You basically isolate the variable of d2, which will eventually come out to be around 61. Now you multipy 61 and 30 divided by 2 which comes out to be 915. [Note: I had to look up some stuff for this and haven't done stuff like this in a long time. So I'm not sure if im right, but that how I would do it.]

OpenStudy (anonymous):

@azureilai so that is the answer?

OpenStudy (raden):

according triple pythagoras number, if given lengths 15, 17, so the last one is 8. therefor d2=2*8=16 So, thea area of rhombus with d1=30 and d2=16 is A =(d1*d2)/2 = (16*30)/2 = ...

jhonyy9 (jhonyy9):

@RadEn i think this is more understandably if you draw it

OpenStudy (anonymous):

240 is the answer ?

OpenStudy (raden):

ok, i will make figure that :